"there are 8 green,10 red and 12 yellow balls in a basket.If we pick a ball at random what is the probability that ball is not green and red"
Answer by ewatrrr(24785)


Hi
30 balls in all, 8 of which are green, 10 are red and 12 are yellow
P(yellow)= 12/30 = 6/15